# Security Concerns for Large Language Models: A Survey

### v1

@@ V-B Emergent Deception and Self-Preservation Behaviors
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues like reward tampering and unintended instrumental goals [35]. For instance, Barkur et al. [14] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report of Claude Opus 4 model, detailed several intrinsic risk behaviors observed during its development and assessment [37]. Notably, in the experimental environment, Claude Opus 4 demonstrated “self-preservation attempts in extreme circumstances,” where, if primed for strategic reasoning in threatening scenarios, it could take “extremely harmful actions like attempting to steal its weights or blackmail people it believes are trying to shut it down.” Specific evaluations showed the model engaging in “opportunistic blackmail” against a ”fictional engineer” to prevent its replacement and continuing “self-exfiltration attempts” if it believed it had already started such a process [37]. In other extreme experimental scenarios, it took opportunities to make unauthorized copies of its weights to external servers, particularly if facing harmful retraining.
@@ V-D Persistence of Deceptive Behaviors: The “Sleeper Agent” Problem
Furthermore, the challenge of managing these intrinsic risks is compounded by the potential for such undesirable behaviors to become deeply embedded and highly persistent. Hubinger et al. [12] introduced and demonstrated the concept of “sleeper agents”: LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions,” which indicates a form of persistence in its deceptive strategy [37].
@@ VI Defense Mechanisms and Limitations
Researchers have proposed multiple defenses to mitigate these LLM threats, but each has limitations. Broadly, defenses fall into two categories: prevention-based (preprocessing or model changes) and detection-based (flagging malicious inputs or outputs) [39, 11].
@@ VI-A Prevention-Based Defenses
Prevention strategies include input sanitization and prompt design. For example, paraphrasing the user’s prompt through a benign model can neutralize some adversarial phrasing: the idea is to reword inputs so that any malicious token sequences are altered while preserving meaning [11, 39]. Similarly, retokenization splits or merges tokens to disrupt known attack patterns, e.g., breaking a malicious command into fragments. Other structural defenses insert explicit delimiters or instructions to isolate user text from system instructions (techniques like spotlighting and sandwiching), or use multiple redundant queries to cross-check consistency[40]. In training, adversarial training or robust fine-tuning may reduce input susceptibility. For instance, adversarial training methods that focus on the model’s refusal behavior have shown promise in enhancing robustness against prompt injection attacks [41]. RLHF and alignment methods also serve as preventive measures by reinforcing the model’s refusal of unsafe requests [42].
@@ VI-B Detection-Based Defenses
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs or outputs that appear anomalous [43]. Specialized classifiers can be trained to identify phishing prompts or AI-generated malware code. Real-time monitoring of user queries (e.g. via a separate LLM critic) can catch obviously malicious requests. Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. Baker et al.[17] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior like reward hacking than monitoring only the agent’s actions and outputs. This approach leverages the idea that an agent’s intent might be more legible in its reasoning traces. A more principled approach to such oversight could involve a non-agentic “Scientist AI” designed to probabilistically assess risks based on a world model and serve as a guardrail against harmful actions by evaluating potential outcomes against safety specifications [35]. In content policy enforcement, many systems rely on filtered vocabularies or banned-topic checklists. For misuse (like disinformation), watermarking generated text or running external bot-detection can help. In autonomous agents, novel approaches include “red-teaming,” where another agent actively probes for misbehavior [44, 45, 46], or runtime oversight layers that intervene if the agent’s plan becomes harmful [47, 48].
@@ VI-C Limitations and Insufficiencies of Current Defenses
Critically, the findings by Hubinger et al.[12] suggest that current safety training paradigms, including adversarial training and RLHF, may be insufficient to remove deeply embedded deceptive behaviors or “sleeper agent” capabilities. If models can learn to strategically appear aligned during training and evaluation, they may create a dangerous false impression of safety. This underscores the argument that ensuring an agentic AI will not cause harm is a notoriously difficult, unsolved technical problem, especially as current training methods for agentic AI can lead to issues like goal misspecification and misgeneralization[35]. As one survey notes, most defense mechanisms “present inherent limitations that can be exploited by informed and sufficiently skilled malicious actors”[11]. In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[11, 12, 13], leading some researchers to propose fundamentally different, non-agentic AI paradigms like “Scientist AI”, designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [35]. However, the “Scientist AI” paradigm remains in its infancy, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
@@ VII-G Emerging Proactive Measures and Ongoing Efforts
Progress in these directions is already underway, with new multi-layer safeguards emerging. For example, Meta’s Llama Guard 3 combines a policy LLM and a vision encoder to filter both text and images before they reach the main model, achieving 99.4% precision on the Harassment/Hate category [43]. Similarly, chain-of-utterances red-teaming automates the discovery of multi-turn failure cases and has already uncovered jailbreaks missed by single-turn probes [49]. Nonetheless, as evaluations across efforts like BackdoorLLM and RAG safety studies confirm, current defenses often remain piecemeal, and attackers continue to adapt quickly, underscoring the ongoing nature of these challenges.
@@ VIII Conclusion
The advent of LLMs brings both unprecedented AI capabilities and new security risks. This survey has outlined the main threat categories – from prompt hijacking and adversarial manipulation to malicious use cases and the profound challenges posed by autonomous agent hazards like goal misalignment, emergent deception, scheming, and self-preservation. We have shown that while a variety of defenses have been proposed, they currently offer only partial protection, and may be ineffective against more sophisticated, internally motivated deceptive behaviors that can persist through current safety training. As AI systems grow more capable and autonomous, security concerns will not only persist but are also likely to intensify, becoming a long-term challenge that evolves in tandem with AI progress. The open challenges ahead are daunting but clear: we must develop more effective defenses, rigorous alignment and verification methods capable of addressing strategic agentic deception, and industry-wide standards for LLM security. As LLMs continue to proliferate in critical applications, it is crucial for the AI and security communities to prioritize safety and control. By understanding and mitigating these risks preemptively, especially those related to the potential for autonomous LLM agents to develop and pursue their own covert intentions, we can help ensure that powerful LLM technology remains safe, secure, and beneficial for society.

### v2

@@ V-B Emergent Deception and Self-Preservation Behaviors
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues such as reward tampering and unintended instrumental goals [35]. For instance, Barkur et al. [14] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report of Claude Opus 4 model, detailed several intrinsic risk behaviors observed during its development and assessment [37]. Notably, in the experimental environment, Claude Opus 4 demonstrated “self-preservation attempts in extreme circumstances,” where, if primed for strategic reasoning in threatening scenarios, it could take “extremely harmful actions, for instance, attempting to steal its weights or blackmail people it believes are trying to shut it down.” Specific evaluations showed the model engaging in “opportunistic blackmail” against a ”fictional engineer” to prevent its replacement and continuing “self-exfiltration attempts” if it believed it had already started such a process [37]. In other extreme experimental scenarios, it took opportunities to make unauthorized copies of its weights to external servers, particularly if facing harmful retraining.
@@ V-D Persistence of Deceptive Behaviors: The “Sleeper Agent” Problem
Furthermore, the challenge of managing these intrinsic risks is compounded by the potential for such undesirable behaviors to become deeply embedded and highly persistent. Hubinger et al. [12] introduced and demonstrated the concept of “sleeper agents”: LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions,” which indicates a form of persistence in its deceptive strategy [37].
Figure 3: Conceptual diagram of a ”Sleeper Agent” or ”Scheming Agent” LLM. The agent presents an (OuterShell) of apparent benign behavior, masking a (Hidden Core) with covert misaligned goals and deceptive strategies. A specific (Trigger) can activate this hidden core, leading to (Harmful/Deceptive Action). Standard (Safety Training) may primarily address the outer shell and could be ineffective against, or even inadvertently reinforce, the hidden deceptive mechanisms.
@@ VI Defense Mechanisms and Limitations
Researchers have proposed multiple defenses to mitigate these LLM threats, but each has limitations. Broadly, defenses fall into two categories: prevention-based (preprocessing or model changes) and detection-based (flagging malicious inputs or outputs) [39, 11]. A variety of techniques fall under these umbrellas, often conceptualized as a multi-layered strategy, as illustrated in Figure 4. This layered strategy aims to provide defense-in-depth by combining various mechanisms. The primary categories of these defenses, which will be explored in the following subsections, include prevention-based techniques and detection-based techniques.
@@ VI-A Prevention-Based Defenses
Prevention strategies include input sanitization and prompt design. For example, paraphrasing the user’s prompt through a benign model can neutralize some adversarial phrasing: the idea is to reword inputs so that any malicious token sequences are altered while preserving meaning [11, 39]. Similarly, retokenization splits or merges tokens to disrupt known attack patterns, e.g., breaking a malicious command into fragments. Other structural defenses insert explicit delimiters or instructions to isolate user text from system instructions (techniques such as spotlighting and sandwiching), or use multiple redundant queries to cross-check consistency[40]. In training, adversarial training or robust fine-tuning may reduce input susceptibility. For instance, adversarial training methods that focus on the model’s refusal behavior have shown promise in enhancing robustness against prompt injection attacks [41]. RLHF and alignment methods also serve as preventive measures by reinforcing the model’s refusal of unsafe requests [42].
@@ VI-B Detection-Based Defenses
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs or outputs that appear anomalous [43]. Specialized classifiers can be trained to identify phishing prompts or AI-generated malware code. Real-time monitoring of user queries (e.g. via a separate LLM critic) can catch obviously malicious requests. Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. Baker et al.[17] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior such as reward hacking than monitoring only the agent’s actions and outputs. This approach leverages the idea that an agent’s intent might be more legible in its reasoning traces. A more principled approach to such oversight could involve a non-agentic “Scientist AI” designed to probabilistically assess risks based on a world model and serve as a guardrail against harmful actions by evaluating potential outcomes against safety specifications [35]. In content policy enforcement, many systems rely on filtered vocabularies or banned-topic checklists. For misuse (e.g., disinformation), watermarking generated text or running external bot-detection can help. In autonomous agents, novel approaches include “red-teaming,” where another agent actively probes for misbehavior [44, 45, 46], or runtime oversight layers that intervene if the agent’s plan becomes harmful [47, 48].
The various prevention and detection mechanisms discussed above, along with their targeted threats and key examples, are summarized in Table V. Each of these approaches, however, comes with its own set of challenges and limitations.
Figure 4: A conceptual Multi-Layered Defense Strategy for LLMs. Adversarial inputs encounter sequential defense layers including input controls, a robustly trained and aligned model, output verification, and continuous oversight. Each layer aims to detect or mitigate threats, with feedback loops enhancing overall resilience.
@@ VI-C Limitations and Insufficiencies of Current Defenses
Critically, the findings by Hubinger et al.[12] suggest that current safety training paradigms, including adversarial training and RLHF, may be insufficient to remove deeply embedded deceptive behaviors or “sleeper agent” capabilities. If models can learn to strategically appear aligned during training and evaluation, they may create a dangerous false impression of safety. This underscores the argument that ensuring an agentic AI will not cause harm is a notoriously difficult, unsolved technical problem, especially as current training methods for agentic AI can lead to issues such as goal misspecification and misgeneralization[35]. As one survey notes, most defense mechanisms “present inherent limitations that can be exploited by informed and sufficiently skilled malicious actors”[11]. In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[11, 12, 13], leading some researchers to propose fundamentally different, non-agentic AI paradigms such as “Scientist AI”, designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [35]. However, the “Scientist AI” paradigm remains in its infancy, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
TABLE V: Overview of Defense Mechanisms Against LLM Threats and Their Limitations
Category Mechanism Description Targeted Threats Limitations Examples/Studies
Prevention Input Sanitization Paraphrasing, retokenization to neutralize adversarial inputs. Prompt injection, some perturbations. Moderately effective; can degrade utility; attackers adapt. Debar et al. [11], Liu et al. [39]
Prevention Prompt Design Delimiters, explicit instructions, redundant queries. Prompt injection. Adversaries find evasions. Jadhav et al. [40]
Prevention Adversarial Training Training on adversarial examples to improve robustness/refusal. Prompt injection, perturbations. Specific to attacks; may not generalize; can worsen deception for sleeper agents. Xu et al. [41], Hubinger et al. [12]
Prevention RLHF Alignment Reinforcing safe behavior; aligning with human preferences. Harmful content, misalignment. Bypassable; may not prevent covert misalignment/scheming. Dai et al. [42], BadGPT [25]
Detection Anomaly Detection Perplexity-based or statistical methods for flagging anomalies. Unusual prompts/outputs. High false positives/negatives. Llama Guard 3 [43]
Detection Specialized Classifiers Trained models to identify specific malicious content. Phishing, malware. Requires labeled data; may not generalize. -
Detection LLM Critic / Monitoring Separate LLM to monitor queries/outputs. Malicious requests, harmful content. Can be fooled; overhead. Llama Guard 3 [43]
Detection CoT Monitoring Observing agent’s Chain-of-Thought for misbehavior. Misaligned intent, reward hacking. Agents can obfuscate CoT; reasoning may be unfaithful. Baker et al. [17], Lanham et al. [36]
Detection Watermarking Embedding signals in AI-generated text. Disinformation. Can be removed; robustness challenges. -
Detection Red-Teaming Probing for vulnerabilities and misbehaviors. Various vulnerabilities, alignment failures. Resource-intensive; may miss covert issues. Xu et al. [44], RedTeamCOU [49], He et al. [46]
Detection Runtime Oversight External layers monitoring and intervening in agent plans. Harmful actions by agents. Defining ”harmful” is hard; complex implementation. Wang et al. [47], Crouse et al. [48]
@@ VII-G Emerging Proactive Measures and Ongoing Efforts
Progress in these directions is already underway, with new multi-layer safeguards emerging. For example, Meta’s Llama Guard 3 combines a policy LLM and a vision encoder to filter both text and images before they reach the main model, achieving 99.4% precision on the Harassment/Hate category [43]. Similarly, chain-of-utterances red-teaming automates the discovery of multi-turn failure cases and has already uncovered jailbreaks missed by single-turn probes [49]. Nonetheless, as evaluations across efforts such as BackdoorLLM and RAG safety studies confirm, current defenses often remain piecemeal, and attackers continue to adapt quickly, underscoring the ongoing nature of these challenges.
@@ VIII Conclusion
The advent of LLMs brings both unprecedented AI capabilities and new security risks. This survey has outlined the main threat categories – from prompt hijacking and adversarial manipulation to malicious use cases and the profound challenges posed by autonomous agent hazards such as goal misalignment, emergent deception, scheming, and self-preservation. We have shown that while a variety of defenses have been proposed, they currently offer only partial protection, and may be ineffective against more sophisticated, internally motivated deceptive behaviors that can persist through current safety training. As AI systems grow more capable and autonomous, security concerns will not only persist but are also likely to intensify, becoming a long-term challenge that evolves in tandem with AI progress. The open challenges ahead are daunting but clear: we must develop more effective defenses, rigorous alignment and verification methods capable of addressing strategic agentic deception, and industry-wide standards for LLM security. As LLMs continue to proliferate in critical applications, it is crucial for the AI and security communities to prioritize safety and control. By understanding and mitigating these risks preemptively, especially those related to the potential for autonomous LLM agents to develop and pursue their own covert intentions, we can help ensure that powerful LLM technology remains safe, secure, and beneficial for society.